OpenStudy (anonymous):

Multiply and simplify:

OpenStudy (anonymous):

The problem is: \[(9+\sqrt{3})^{2}\]

OpenStudy (lilg132):

\[84 + 18\sqrt{?}\]

OpenStudy (anonymous):

The answer choices are: A. \[81+2\sqrt{3}\] B. \[84+18\sqrt{3}\] C. 84

OpenStudy (lilg132):

\[84 + 18\sqrt{3}\]

jimthompson5910 (jim_thompson5910):

\[(9+\sqrt{3})^{2}\] \[(9+\sqrt{3})(9+\sqrt{3})\] \[(9)^2+2*9*\sqrt{3}+(\sqrt{3})^2\] \[81+18*\sqrt{3}+3\] \[84+18*\sqrt{3}]

jimthompson5910 (jim_thompson5910):

So choice B